calculate percentage error  in determination of time period of a pendilum , 


 


T = 2π √l/g


where l and g are measured with ±1% and ±2%

Grade:11

3 Answers

Nimish Singh
11 years ago

     dt/t*100 = 1/2 *dl/l*100 + 1/2 *dg/g*100

          dt/t*100= 1/2*1+1/2*2=1/2+1=3/2=1.5%

        Therefore percentage error is ±1.5%

Please like my answer if it is correct

tejas naresh patel
11 years ago

no idea

Abhishekh kumar sharma
10 years ago

 dt/t*100 = 1/2 *dl/l*100 + 1/2 *dg/g*100

          dt/t*100= 1/2*1+1/2*2=1/2+1=3/2=1.5%

      percentage error is ±1.5%